Project K costs $55,000, its expected cash inflows are $15,000 per year for 11 years, and its WACC is 9%. What is the project's payback? Round your answer to two decimal places.
Answer #1

Payback period is the time taken for the cumulative cash flows to equal zero

Payback period = 3 + (cash flow required in year 4 for cumulative cash flows to equal zero / year 4 cash flow)

Payback period = 3 + ($10,000 / $15,000) = 3.67 years

IH Bc Cumulative Cash Flow Cash Flow 0 ($55,000) ($55,000) $15,000 ($40,000) 2 $15,000 ($25,000) 3 $15,000 ($10,000) 4 $15,00